OpenStudy (sissyedgar):

What is the solution to the system of equations represented by the two equations? y = 2/3x y = -2/3x + 4 ( , )

OpenStudy (kainui):

I would suggest adding both equations together

pooja195 (pooja195):

2/3=-2/3x+4

OpenStudy (kainui):

Almost, but not quite.

pooja195 (pooja195):

@ucreategames can you show how you got that?

OpenStudy (kainui):

y = 2/3x y = -2/3x + 4 add them: y+y = 2/3x -2/3x+4 2y=4 y=2

OpenStudy (sissyedgar):

SO how would I out that in the ( , )? {4,2} @Kainui

OpenStudy (kainui):

Well as it is I've only shown you how to solve for y, you still need to solve for x. Once you do that, then you put them in like this, replacing the letters with their values: (x,y)

OpenStudy (sissyedgar):

ooo, ok (x,2)

OpenStudy (sissyedgar):

How do I solve for x @Kainui

OpenStudy (sissyedgar):

...